# What constitutes a Deep Fake? The blurry line between legitimate processing and manipulation under the EU AI Act

### v1

@@ 3.3 Summary of main critical legislative issues in Article 50 AI act
For the later discussion of the challenges, we would like to summarise our main critique relating to the legislative technique w. r. t. the transparency obligation for deep fakes and synthetic content.
@@
First, for the definition of a deep fake in Article 3(60) which is used in Article 50, it is unclear how the manipulation of content in deep fakes should be understood. This is particularly true for images, as assumptions and settings are already made by the photographer during the capture of the image, meaning that reality is never truly represented.
@@
Second, Article 50(2) contains exceptions for assistive functions for standard editing. What can be considered as such a function? How broad is the definition of standard editing? Can the Google “best take” function, which is a newly developed feature, be considered as standard editing?
@@
Third, Article 50(2) also contains an exception for not substantially altering the content of an image. How broad is the definition of altering the content? What can be considered content in the sense of Article 50(2) in cases of digital audio, images, and videos? For example, in the case of images, do we need to consider the pixel-space or the visible space perceptible by humans? Substantive manipulations in the pixel space might not change human perception, and on the other hand, small perturbations in the pixel space can change the perception dramatically.
@@
4 Challenges and edge cases
@@
4.1 Differentiation between legitimate AI-based image processing and deep fake generation
@@
As discussed above, an image never depicts the objective truth. Traditional processing steps (see section 2.1) are commonly considered to lead to authentic outcomes. Moon photos processed by Samsung smartphones (see section 2.2) do not depict the reality as recorded by an image sensor, but are closer to the human perception of the moon than unaltered photos. As a result, a moon photo processed in this way does not constitute a deep fake per se; as the appearance of the moon is not expected to change, the image can in fact be considered as authentic. In summary, the photos depict the actual moon, and the AI function merely helps to overcome a technical limitation; the moon is therefore authentic in the resulting photos.
@@
It is, however, conceivable that an image containing other objects besides the (post-processed) moon might have to be considered as a deep fake: A viewer might assume that those other objects are depicted with the same quality as the moon, and thus falsely consider the lack of structure in these objects as authentic. A deep fake would also be created if the AI system falsely identifies an object as the moon; for example, Reddit user ibreakphotos (see section 2.2) demonstrated Samsung’s moon shot feature by taking photos of a computer screen, not the actual moon.
4.2 Best Take
@@
While replacing an image of the moon with a better version may only result in deep fakes in edge cases, Google’s “best take” seems to be an obvious example for manipulation: By exchanging a face in a group photo, the feature creates a depiction of a situation that never occured. For such an image to be considered a deep fake, it would have to “falsely appear to a person to be authentic or truthful” (Article 3 (60)). If the photo is objectively authentic and truthful, it is not a deep fake. On one hand, in a typical group photo setting, an average viewer would probably still consider the resulting photo as authentic. The smile which is inserted existed within a couple of seconds from the remaining photo being taken. On the other hand, the ten second time frame of the best take feature is sufficient for a mood change. A person might have stopped smiling while the rest of the group laughs about a joke at their expense.
@@
As a consequence, we assume that such a group photo may well constitute a deep fake. The feature might still be “an assistive function for standard editing” according to Article 50(2), in which case the transparency obligation for providers of AI systems would not apply. Article 50(4), which applies to deployers of AI sytems, does not include a similar exception; we still discuss whether features similar to “best take” can be considered as an assistive function for standard editing.
@@
One could argue that combining parts of several images goes beyond editing an image. The wording of the law, however, does not specifically refer to editing of only one individual image. Combining information from several images has also been common at least since the advent of digital photography. The dynamic range of a picture can be improved this way (HDR composites). In case of panoramic shots, the individual pictures are often taken several seconds apart, but are not usually considered as problematic. Both HDR and panoramic photos are common enough to be considered as standard editing. The use of AI alone does not suffice to not consider a post-processing tool as a standard editing tool—otherwise, the exeption in Article 50(2) would not make sense.
@@
We conclude that there is indeed a potential for the generation of deep fakes, which do not fall under the exception of Article 50(2).
4.3 Altering an image: Pixel space vs. visible space?
Besides the “standard editing” exception, the transparency requirement of Article 50(2) does not apply to AI systems that “do not substantially alter the input data provided by the deployer or the semantics thereof”. We consider this wording to be problematic. While changes to the semantics are the exact reason why legislators deemed it necessary to regulate deep fakes, an exception for minor changes to the input data introduces an unnecessary backdoor. Since the exception for minor changes to the input is given equal weight alongside the semantics exception, there is little scope for interpreting the term ‘input data’ to include content-related aspects. The extent of alteration of input data is not an appropriate measure for the impact of that alteration. Fig. 4 shows an original image and two edited versions. In both modifications, the average alteration of the input data (measured by the root mean squared differences of the red, green, and blue channels for each pixel) w.r.t. the original image is identical, but the semantics are clearly different.
The exception of Article 50(2) applies when either the input data are not substantially altered, or the semantics are not substantially altered. For further discussion we use the stylized example in Figure 4. In Figure 4(b), the semantics are substantially altered compared to Figure 4(a). But are the input data substantially altered? The manipulation in Figure 4(c) does not lead to a significant perceptible visual difference. The root mean squared differences of Figures 4(b) and Figure4(c) compared to Figure 4(a) are identical. Thus, one can argue that even in the case of Figure 4(b), the input data are not substantially altered. Consequently, the exception applies and Figure 4(b) would fall under the scope of the exception of Article 50(2). Article 50(2) would not be applicable in the stylized example.
@@
Figure 4: Difference in the pixel space vs. visible space. (a) Original image. (b) Added a pistol to the image. (c) Image with slightly modified intensity around the hand. Figure (b) and Figure (c) have the same root mean squared difference (1.83 pixel) compared to Figure (a). To reduce potential harm and misuse we used a stylized pistol in case of Figure (b). Pistol icon from Flaticon.com.
5 Conclusion and Outlook
@@
The AI Act’s definition of deep fakes, and the obligations applying to providers and deployers of AI systems that generate deep fakes, are less clear than they appear at first glance. As (digital) photos are never shown exactly as recorded by the image sensor, there has always been room for interpretation and alteration of their visual impact. The delineation between legitimate processing (with the result of an image that is still considered as “authentic”) and deep fakes (which ´´would falsely appear to a person to be authentic or truthful”) is a problematic one, and the mere use of AI tools in image processing does not suffice to cross the threshold of a deep fake. We note that regulating the use of AI systems for image manipulation does make sense, as those tools are easily accessible and produce convincing images. Moreover, even legitimate AI-based image manipulation tools used without malicious intent introduce risks, as mistakes can happen in rare cases. Considering the reliance on photos (e.g., in journalism or as evidence in court), transparency requirements should not be interpreted too narrowly. A stronger focus on human oversight might have been a more appropriate approach for regulating the use of AI in image editing.
@@
We expect discussions around deep fakes in other law areas as well. For example, there is an increasing use of deep fakes in pornography, including CSAM (child sexual abuse material) and youth pornography. Prosecutors face challenges of identifying such material and have to consider its consequences for prosecution, and legislators may be urged to react to this trend as well.
@@
6 Acknowledgments
We would like to thank Nils Wiedemann for helpful comments on the manuscript. In addition, the authors would like to thank Thomas Wagner for providing images of the moon. Finally, we are grateful to the anonymous reviewers of the ACM Symposium on Computer Science and Law.

### v2

@@ 3.3. Summary of main critical legislative issues in Article 50 AI Act
For the later discussion of the challenges, we would like to summarize our main critique relating to the legislative technique w. r. t. the transparency obligation for deep fakes and synthetic content.
@@
First, for the definition of a deep fake in Article 3(60) AIA that is used in Article 50 AIA, it is unclear how the manipulation of content in deep fakes should be understood. This is particularly true for images, as assumptions and settings are already made by the photographer during the capture of the image, meaning that reality is never truly represented.
@@
Second, Article 50(2) AIA contains exceptions for assistive functions for standard editing. What can be considered as such a function? How broad is the definition of standard editing? Can the Google “best take” function, which is a newly developed feature, be considered as standard editing?
@@
Third, Article 50(2) AIA also contains an exception for not substantially altering the content of an image. How broad is the definition of altering the content? What can be considered content in the sense of Article 50(2) AIA in cases of digital audio, images, and videos? For example, in the case of images, do we need to consider the pixel space or the visible space perceptible by humans? Substantive manipulations in the pixel space might not change human perception, and on the other hand, small perturbations in the pixel space can change the perception dramatically.
@@
4. Challenges and edge cases
@@
4.1. Differentiation between legitimate AI-based image processing and deep fake generation
@@
As discussed above, an image never depicts the objective truth. Traditional processing steps (see Section 2.1) are commonly considered to lead to authentic outcomes. Moon photos processed by Samsung smartphones (see Section 2.2) do not depict reality as recorded by an image sensor but are closer to the human perception of the moon than unaltered photos. As a result, a moon photo processed in this way does not constitute a deep fake per se; as the appearance of the moon is not expected to change, the image can, in fact, be considered authentic. In summary, the photos depict the actual moon, and the AI function merely helps to overcome a technical limitation; the moon is, therefore, authentic in the resulting photos.
@@
It is, however, conceivable that an image containing other objects besides the (post-processed) moon might have to be considered a deep fake: A viewer might assume that those other objects are depicted with the same quality as the moon, and thus falsely consider the lack of structure in these objects authentic. A deep fake would also be created if the AI system falsely identifies an object as the moon; for example, Reddit user ibreakphotos (see Section 2.2) demonstrated Samsung’s moon shot feature by taking photos of a computer screen, not the actual moon.
4.2. Best Take
@@
While replacing an image of the moon with a better version may only result in deep fakes in edge cases, Google’s “best take” seems to be an obvious example of manipulation: By exchanging a face in a group photo, the feature creates a depiction of a situation that never occurred. For such an image to be considered a deep fake, it would have to “falsely appear to a person to be authentic or truthful” (Article 3(60) AIA). If the photo is objectively authentic and truthful, it is not a deep fake. On one hand, in a typical group photo setting, an average viewer would probably still consider the resulting photo as authentic. The smile that is inserted existed within a couple of seconds of the remaining photo being taken. On the other hand, the ten-second time frame of the Best Take feature is sufficient for a mood change. A person might have stopped smiling while the rest of the group laughs about a joke at their expense.
@@
As a consequence, we assume that such a group photo may well constitute a deep fake. The feature might still be “an assistive function for standard editing” according to Article 50(2) AIA, in which case the transparency obligation for providers of AI systems would not apply. Article 50(4) AIA, which applies to deployers of AI systems, does not include a similar exception; we still discuss whether features similar to “best take” can be considered an assistive function for standard editing.
@@
One could argue that combining parts of several images goes beyond editing an image. The wording of the law, however, does not specifically refer to the editing of only one individual image. Combining information from several images has also been common, at least since the advent of digital photography. The dynamic range of a picture can be improved this way (HDR composites). In the case of panoramic shots, the individual pictures are often taken several seconds apart. This is not usually seen as problematic. Both HDR and panoramic photos are common enough to be considered standard editing. The use of AI alone does not suffice to not consider a post-processing tool as a standard editing tool—otherwise, the exception in Article 50(2) AIA would not make sense.
@@
We conclude that there is indeed a potential for the generation of deep fakes that do not fall under the exception of Article 50(2) AIA.
4.3. Altering an image: Pixel space vs. visible space?
Besides the “standard editing” exception, the transparency requirement of Article 50(2) AIA does not apply to AI systems that “do not substantially alter the input data provided by the deployer or the semantics thereof”. We consider this wording to be problematic. While changes to the semantics are the exact reason why legislators deemed it necessary to regulate deep fakes, an exception for minor changes to the input data introduces an unnecessary backdoor. Since the exception for minor changes to the input is given equal weight alongside the semantics exception, there is little scope for interpreting the term ‘input data’ to include content-related aspects. The extent of alteration of input data is not an appropriate measure for the impact of that alteration. Fig. 4 shows an original image and two edited versions. In both modifications, the average alteration of the input data (measured by the root mean squared differences of the red, green, and blue channels for each pixel) w.r.t. the original image is identical, but the semantics are clearly different.
The exception of Article 50(2) AIA applies when either the input data (provided by the deployer) or their semantics are not substantially altered. For further discussion, we use the stylized example in Figure 4. In Figure 4(b), the semantics are substantially altered compared to Figure 4(a). But are the input data substantially altered? The manipulation in Figure 4(c) does not lead to a significant perceptible visual difference. The root mean squared differences of Figures 4(b) and Figure4(c) compared to Figure 4(a) are identical. Thus, one can argue that even in the case of Figure 4(b), the input data are not substantially altered. Consequently, the exception applies and Figure 4(b) would fall under the scope of the exception of Article 50(2) AIA. Article 50(2) AIA would not be applicable in the stylized example.
@@
[…]
An image of a man in front of a bright background. The man seems to be pointing at something with his right index finger.
@@
[…]
An image of a man in front of a bright background. The image is identical to the leftmost one, with the exception of a cartoon-like drawing of a pistol placed on top of the right hand.
@@
[…]
An image of a man in front of a bright background. The man seems to be pointing at something with his right index finger. The image appears identical to the leftmost one.
@@
Figure 4. Difference in the pixel space vs. visible space. (a) Original image. (b) Added a pistol to the image. To reduce potential harm and misuse, we used a stylized pistol. (c) Image with slightly modified intensity around the hand. Figure (b) and Figure (c) have the same root mean squared difference (1.83 pixel) compared to Figure (a). Pistol icon designed by freepik from Flaticon (flaticon.com).
5. Recommendations
Several issues within the AI Act and the corresponding obligations have been raised above. In this section, we want to highlight possible recommendations to mitigate these issues.
Regarding the definition of deep fakes in Article 3(60) AIA, we have demonstrated that distinguishing manipulated content from legitimate processing is challenging. However, the wording of the law gives the competent authorities and courts enough leeway to find sensible solutions for individual cases. Similarly, while examples such as the “best take” feature show that while a clarification of the term “standard editing” by the legislator might be helpful, the issue can be resolved by the judiciary. As discussed above, the term should not be applied too broadly. Even a seemingly harmless feature like “best take” has the potential to generate deep fakes under specific circumstances.
The transparency requirements (laid down in Article 50 AIA) on the other hand, should be clarified by the legislator. From our point of view, the exception of “not substantially altered input data” in Article 50(2) AIA is problematic, as alteration of data (as the term “data” is generally understood) is not meaningful for human perception (as we show in Figure 4). The semantics of an image can be changed in a harmful way, even if there is no substantial alteration of input data.
The exception should therefore be removed, as the other exception in the same sentence, referring to substantially altered semantics, already covers the relevant case of a changed meaning.
Furthermore, we argue that the unclear relationship between the obligations of deployers and providers in Article 50(2) AIA and Article 50(4) AIA leads to unnecessary confusion. It remains unclear whether the deployer and the provider must simultaneously label the output data. This becomes even more pressing as Article 50(2) AIA targets deployers and synthetic content, while Article 50(4) AIA targets providers and deep fakes. The AI Act would benefit from an explicit and congruent statement of the relationship.
As the AI Act is not likely to be amended in this way in the near future, the competent authorities and judiciary will have to deal with its shortcomings in the meantime. In addition, the European Artificial Intelligence Board, whose task is (among others) to “collect and share technical and regulatory expertise and best practices among Member States” (Article 66(b) AIA), may contribute to a better and harmonized understanding.
By making core issues of the regulation of deep fakes explicit, the paper at hand aims to support the competent bodies in finding pragmatic solutions in line with the legislator’s intent.
6. Conclusion and Outlook
@@
The AI Act’s definition of deep fakes, and the obligations applying to providers and deployers of AI systems that generate deep fakes, are less clear than they appear at first glance. As (digital) photos are never shown exactly as recorded by the image sensor, there has always been room for interpretation and alteration of their visual impact—even before the widespread use of AI. The delineation between legitimate processing (with the result of an image that is still considered as “authentic”) and deep fakes (which “would falsely appear to a person to be authentic or truthful”) is a problematic one, and the mere use of AI tools in image processing does not suffice to cross the threshold of a deep fake. We note that regulating the use of AI systems for image manipulation is indeed crucial, as those tools are easily accessible and produce convincing images. Moreover, even legitimate AI-based image manipulation tools used without malicious intent introduce risks, as mistakes can happen in rare cases. Considering the reliance on photos (e.g., in journalism or as evidence in court), transparency requirements should not be interpreted too narrowly. A stronger focus on human oversight might be a more appropriate approach for regulating the use of AI in image editing.
@@
We expect discussions around deep fakes in other law areas as well. For example, there is increasing use of deep fakes in pornography, including CSAM (child sexual abuse material) and youth pornography. Prosecutors face challenges of identifying such material and have to consider its consequences for prosecution, and legislators may be urged to react to this trend as well.
@@
7. Acknowledgments
We would like to thank Nils Wiedemann and Roman Wetenkamp for their helpful comments on the manuscript. In addition, the authors would like to thank Thomas Wagner for providing images of the moon. Finally, we are grateful to the anonymous reviewers of the ACM Symposium on Computer Science and Law.